Solving Systems Using Elimination. Your math test has 38 questions and is worth 200 points. The test consist of multiple choice questions worth 4 points each and open ended questions worth 20 points each. How many of each type of question are there?
So far I have
x + y = 38
4x +20y = 200

Answer by ikleyn(52788)

.

Your start is fine.


The next steps are

     x +  y =   38    (1)
    4x + 20y = 200    (2)


To solve by Elimination, multiply equation (1) by 4 (both sides). Keep equation (2) as is.


    4x + 4y =  4*38   (1')
    4x + 20y = 200    (2')


Now subtract equation (1') from equation (2').  The terms " 4x " will kill each other, and you will get single equation 
for variable "y" only (!)


         20y - 4y = 200 - 4*38 

         16y      = 48

           y      = 48/16 = 3.


It how the Elimination method works.


Finally, substitute the value y= 3 into equation (1) to get "x".


          x + 3 = 38

          x     = 38-3 = 35.


ANSWER.  x= 35;  y= 3.

Solved.

--------------------

In this site, there is a large group of lessons on solving 2x2-system of linear equations
    - Solution of the linear system of two equations in two unknowns by the Substitution method
    - Solution of the linear system of two equations in two unknowns by the Elimination method
    - Solution of the linear system of two equations in two unknowns using determinant
    - Geometric interpretation of the linear system of two equations in two unknowns
    - Useful tricks when solving systems of 2 equations in 2 unknowns by the Substitution method
    - Solving word problems using linear systems of two equations in two unknowns

    - Oranges and grapefruits
    - Using systems of equations to solve problems on tickets
    - Three methods for solving standard (typical) problems on tickets
    - Using systems of equations to solve problems on shares
    - Using systems of equations to solve problems on investment
    - Two mechanics work on a car
    - The Robinson family and the Sanders family each used their sprinklers last summer
    - Roses and vilolets
    - Counting calories and grams of fat in combined food
    - A theater group made appearances in two cities
    - HOW TO algebreze and solve this problem on 2 equations in 2 unknowns
    - One unusual problem to solve using system of two equations
    - Solving mentally word problems on two equations in two unknowns
    - Solving systems of non-linear equations by reducing to linear ones
    - System of equations helps to solve a problem for the Thanksgiving day
    - Using system of two equations to solve the problem for the day of April, 1

    - OVERVIEW of lessons on solving systems of two linear equations in two unknowns

Also,  you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Systems of two linear equations in two unknowns".


Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.